OpenStudy (anonymous):

How many distinguishable permutations can be made using "SEATTLE". Please explain how to do it because I need to know how to do these. PLEAS TELL ME HOW TO GET 1260 BECAUSE THAT"S SUPPOSED TO BE THE RIGHT ANSWER!

OpenStudy (lgbasallote):

how many letters does SEATTLE have?

OpenStudy (anonymous):

7 letters

OpenStudy (lgbasallote):

how many times did E repeat?

OpenStudy (anonymous):

twice

OpenStudy (lgbasallote):

how many times did T repeat?

OpenStudy (anonymous):

twice

OpenStudy (lgbasallote):

so the formula would be \[\large \frac{7!}{2!2!}\]

OpenStudy (lgbasallote):

7 is the number of letters..that becomes the numerator...two letters repeated..both repeated twice so they go in the denominator

OpenStudy (anonymous):

Is that because n!/(n - r)! = 7!/0! = 7!. And then you divide by the number of repeated letters? 2 letters were repeated twice, so you divide by 2!2!?

OpenStudy (lgbasallote):

purdy much..for example LGBASALLOTE has 11 letters..l repeats thrice...a repeats twice..so the formula for it is \[\frac{11!}{3!2!}\]

OpenStudy (anonymous):

Alright thanks!